OpenStudy (anonymous):

What are the intercepts of y = 4x − 2?

OpenStudy (anonymous):

x-intercept: (12,0) y-intercept: (0,−2)

OpenStudy (mathstudent55):

Since the equation is in the y = mx + b form, you can read directly b, the y-coordinate of the y-intercept. Since b = -2, the y-intercept is (0, -2). To find the x-intercept, let y = 0, and solve for x.

OpenStudy (anonymous):

Notice how y=4x%2B2 is in slope-intercept form y=mx%2Bb where m is the slope and b is the y-intercept. So this shows us that the slope is m=4 and the y-intercept is b=2 which is the point (0,2)

OpenStudy (anonymous):

Equations in this form are easy to graph, since the slope of line is m, and the y-intercept of the line (the point where it crosses the y-axis) is b. Since the equation is given in slope-intercept form, we know immediately that the line crosses the y-axis at (0, 3) and has slope –2.

OpenStudy (mathstudent55):

y = 4x - 2 Let y = 0: 0 = 4x - 2 4x = 2 x = 1/2 The x-intercpet is: (1/2, 0)

OpenStudy (nincompoop):

do you still need help @CAPTREDEYES420